A widget assembly process was timed for 75 cycles and had an average of 1.5 minutes per piece. The performance rating was 95%, and workday allowances are 10 percent. Determine the following: a) Observed time b) Normal time c) Standard time d) How might this information be used?
Answer #1

A) observed time

Average time = 1.5 minutes

B) normal time

1.5 minutes *. 95 = 1.425 minutes

C) standard time

AF = 1/(1-. 10) = 1.111

1.425 * 1.111 = 1.58 minutes

D) this information is used in industrial engineering, the standard time is the time required by an average skilled operator, working at a normal pace, to perform a specified task using a prescribed method. observed time. the time it takes an operator to perform a task when being observed for time study purposes.The time required by a trained worker to perform a task at a normal pace. The total of all the normal elemental times constituting a cycle or operation.